OpenStudy (anonymous):

Find the Domain. x-9/x^2-81

OpenStudy (anonymous):

x-9 x-9 1 --------- = ------------ = ------- (x)² - (9)² (x+9) (x-9) x+9 now u can do the rest.....................

jhonyy9 (jhonyy9):

x-9 1 --------- = ------ solutions are all numbers without -9 than the denominator is (x-9)(x+9) x+9 zero

myininaya (myininaya):

domain is all real numbers except x=-9 and x=9

OpenStudy (anonymous):

thats what i thought too but it keeps saying its wrong on my hw

myininaya (myininaya):

you have a hole in the function at x=9 (so there is discontinuity here) you have that the function has a vertical asymptote at x=-9(so there is discontinuity here) the function is continuous on its domain since the function is not continuous at x=9 and x=-9 then these numbers are not elements of the domain

myininaya (myininaya):

all real numbers but x=9 and x=-9 is the answer if it is saying it is wrong then it is wrong unless you function is not \[f(x)=\frac{x-9}{x^2-81}\]

OpenStudy (anonymous):

oh nvm.. it wanted me to not check all real numbers.. which is weird bc i dont know why it wouldnt want me to include that in the answer? any ideas?

myininaya (myininaya):

what does your question exactly say?

OpenStudy (anonymous):

It says: Find the Domain. then it gave me the function then the directions were select all that apply.. & i had originally checked in [all real number, x does not equal -9 and x does not equal 9] but when i submitted it it said wrong so i tried it without all real numbers and it was correct.. weird

myininaya (myininaya):

that is weird so there were 2 options one saying all real numbers, x doesn't =-9 and x doesn't =9 and another saying x doesn't=-9 and x doesn't =9 ?

myininaya (myininaya):

but it said you could select more than one right?

myininaya (myininaya):

you should select both then and i don't know what the other choices are but i know you should select both of those

OpenStudy (anonymous):

no.. for example: all real numbers would be a option, x does not equal 9 would be a option and so forth. and you would check the box next to them

OpenStudy (anonymous):

so it was correct when i excluded all real numbers and im not really sure why it would have me do that

myininaya (myininaya):

well it isn't all real numbers because we are excluding 9 and -9

myininaya (myininaya):

and 9 and -9 are real numbers

OpenStudy (anonymous):

ohhh that makes sense!

myininaya (myininaya):

the whole set of real numbers is not domain the real numbers - x=-9 and x=9 is the domain

OpenStudy (anonymous):

oh ok , got cha! thanks! its so simple lol

myininaya (myininaya):

yes but that is kindof confusing i suppose its like they want to trick you

myininaya (myininaya):

lol

OpenStudy (anonymous):

haha yeah it seemed that way. thank you.

myininaya (myininaya):

np